OpenStudy (anonymous):

for what values of a and b does the function f(x)= x^3 + ax^2 + bx + 2 have a local maximum when x = -3 and a local minimum when x = -1?

sam (.sam.):

find f '(x)

OpenStudy (anonymous):

after that?

OpenStudy (anonymous):

f(x)=x^3 + ax^2 + bx + 2 f'(x)=2x^2 +2ax+b Isn't f'(x) = 3x^2 +2ax +b. 3 instead of 2 when differentiating.

sam (.sam.):

oh man

OpenStudy (anonymous):

tiaph has a point.

OpenStudy (anonymous):

Think you need to find f''(x) also to differentiate the values of x that is the minimum and the maximum point, if not you will have 2 set of a and b values.

sam (.sam.):

my monitor is far i can't focus

sam (.sam.):

f(x)=x^3 + ax^2 + bx + 2 f'(x)=3x^2 +2ax+b --------------------------- From the given points, x=-3 x=-1 x+3=0 x+1=0 (x+3)(x+1)=0 x^2+4x+3=0 multiply whole equation by 3, 3x^2+12x+9=0 Compare 3x^2 +2ax+b 3x^2+12x+9 2a=12 b=9

sam (.sam.):

there we go